Shift summary for owl shift 7/15/99

Robert Tschirhart (tsch@fnal.gov)
Thu, 15 Jul 1999 06:51:40 -0500

Shift Summary for owl shift, 07/15/99

On Shift: Pat Toale, Bob Tschirhart

Hours of beam available at all intensities: 6h 15min
Hours of beam available at nominal intensity: 6h 15min
Hours of E832 data (special runs not included): 6h 05min
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Beam losses in Main Injector, minor DAQ probs.

========================================================================
Runs ended this shift:

Run 13911 - BEAM_832 (443K events) tape change.

Tapes brought to FCC this shift: QK*131-133, QKB045
========================================================================

Accesses during shift: none
========================================================================

Overview of shift: 832 data taking at ~6.8 E12

Smooth shift, occasional ktevana errors, HCC inefficiency warnings.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13911 00:43 Tapes are full. This run stops. We load a new
tape set, QK*134-136, QKB046.

13912 00:49 Fresh start, start beam832 run 13912 with updates
to the banana monitoring code.

13912 00:10 Intensity exceeds 7.1E12, we call MCR for trim.

13912 00:45 Onsplit mount request for QKNL11 serviced.

13912 02:45 Ghost message from pipeline frequently occurs,
fixed later with a reboot of pipeline crate #03.

13912 04:00 New Banana maps functioning properly with new
time offsets. Monitoring is still messed up,
and will be updated soon with a fix.

13912 05:15 CsI channel 3114 has an unusually large E/P (1.08)
not due to low electron statistics. Laser data
for this channel looks ok. This channel should
continue to be monitored.

13912 05:30 Onsplit mount request for QKCC06 serviced.

13912 06:10 Getting DAQ deadtime of 0.5->2.0 seconds/spill
sometimes. Not associated with high intensity
spills, not clear what is causing it.

13912 06:15 Investigating beam losses in the Main Injector,
No Estimate.

13912 06:20 Pipeline crate #03 rebooted. This seems to have
fixed ghost status messages from pipeline system.

========================================================================
Follow up Notes on Previous Shift

We followed the orders of not flipping the analysis magnet field and
not doing a muon run.

========================================================================
Items to Follow up on Next Shift

Keep an eye on the E/P (csigain) of CsI 3114.